GlueX TOF Meeting, July 28, 2015

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, July 28, 2015
11:00 am EDT
JLab: CEBAF Center, Room F326/327


  1. Announcements
  2. Review of minutes from July 14
  3. Calibration Status FSU
  4. Calibration Status JLab



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at .



  • FSU: Hussein Al Ghoul, Paul Eugenio, Sasha Ostrovidov, Aristeidis Tsaris
  • JLab: Brad Cannon, Mark Ito (chair), Beni Zihlmann

Review of minutes from July 14

We looked at the minutes.

No movement on the CAEN TDC bin calibration yet. Beni will talk to Sergey Furletov about installing Bryan Moffit library into the online system.

Calibration Status FSU

  1. Sasha trying to use reflections to get at time-walk corrections. He tried looking at pulses with two samples, but usually there are not TDC hits for the small reflections.
  2. Tried taking cosmic data, but the IOC's were down; he could not turn on the HV. Beni told us you have to re-start the IOC's from the near-top-level GUI. He will have a look at the system after the meeting.
  3. There was trouble starting the DAQ from gluon47. None of the windows came up. We suggested that Sasha consult with Sergey Furletov before attempting to start up the DAQ remotely. There may be conditions in the Hall that prevent doing so.

Calibration Status JLab

Beni led us through his wiki page on light attenuation measurements from data. Please see the page for details and figures. In contrast to previous work, he tried fitting to a single exponential to facilitate comparison with the FSU bench results which also used a single exponential. He finds a mean attenuation length of 225 cm with an RMS spread of about 20 cm. This does not agree very well with the FSU results.

Sasha pointed out that the FSU measurements were made with a UV laser and were done with un-wrapped scintillator. Meaningful comparison may be difficult.

Calibration Training

Beni met with Brad yesterday. Brad will be at the lab for the next two weeks. They succeeded in building Beni's code in Brad's account and running the jobs on run 3180. Brad is going through the code now to understand the system.

He will subscribe to the PID and Calibration mailing lists and will attend the calibration meetings as well.